OpenStudy (anonymous):

A company’s network has computers. All communications within the network have to be encrypted. Assuming each pair of computers needs one key for encryption, how many keys will have to be distributed by a company’s Key Distribution Center if all keys have to be different?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

if there are N computers, then \[\sum_{k=1}^{N-1} k \] keys need to be distributed

OpenStudy (anonymous):

ianhardinger, did you understand the answer?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

i'm trying to make sense of it by looking through my notes- in a computing class that went from calculus to statistics in one lesson...I like calculus....

OpenStudy (anonymous):

suppose there are 3 computers, A,B and C. you'd need one key for A-B one key for B-C one key for A-C so, for three computers, the answer is 3 keys. suppose you have four computers A,B,C,D you need a key for A-B,A-C,A-D B-C,B-D C-D so totally 7 keys.

OpenStudy (anonymous):

for N=3, \[\sum_{k=1}^{N-1}k = \sum_{k=1}^{2}k = 1+2 =3\] for N=4 \[\sum_{k=1}^{N-1}k = \sum_{k=1}^{3}k = 1+2+3 =6\]

OpenStudy (anonymous):

I understand when you state it like that

OpenStudy (anonymous):

sorry, I meant 6 keys in the previous post. A-B,A-C,A-D B-C B-D C-D totally 6 keys

OpenStudy (anonymous):

Thanks for your help. This is a pretty nifty little site.

OpenStudy (anonymous):

you are welcome. mark this thread as has good answer so that others don't waste time coming here and help

OpenStudy (anonymous):

where do i say that

OpenStudy (anonymous):

click on the goog answer button next to any reply.

OpenStudy (anonymous):

* Good Answer*

OpenStudy (anonymous):

it doesn't say that anywhere...I'm having JAVA issues...

OpenStudy (anonymous):

oh well. thats alright